A system with effective mass (m) has a potential energy given by ;-
u(x) = Uo (-2 (x/xo)² + (x/xo)⁴)
(b) find the points where the force on the particle is zero
(e) classify these points as stable or unstable.
(f) calculate the value of U(x)/Uo at these equilibrium points.
1
Expert's answer
2020-09-01T07:37:11-0400

"U(x) = U_o (-2 (\frac{x}{x_o})^2 + (\frac{x}{x_o})^4)"

"(b)" For force on the particle to be zero, "F=-\frac{dU}{dx}=0"

"-\frac{dU}{dx}=-U_o[-4(\frac{x}{x_o})(\frac{1}{x_o})+4(\frac{x}{x_o})^3(\frac{1}{x_o})]=0\implies" "(\frac{x}{x_o})(1-(\frac{x}{x_o})^2)=0"

( As "U_o,x_o,4" are constants)

"\implies \frac{x}{x_o}=0\implies x=0" OR

"\implies (\frac{x}{x_o})^2=1 \implies x=+x_o" OR "x=-x_o" .

"(e)" To determine stable or unstable equilibrium, we check "-\frac{d^2U}{dx^2}" .

If it is greater than zero,it is unstable equilibrium.

If it is lesser than zero,it is stable equilibrium.

If it is zero,it is neutral equilibrium.

"-\frac{d^2U}{dx^2}=-U_o[-\frac{4}{x_o^2}+12(\frac{x}{x_o})^2(\frac{1}{x_o})^2]"

For "x=0," "-\frac{d^2U}{dx^2}=-U_o[-\frac{4}{x_o^2}+12(\frac{x}{x_o})^2(\frac{1}{x_o})^2]=" "-U_o[-\frac{4}{x_o^2}+12(\frac{0}{x_o})^2(\frac{1}{x_o})^2]=\frac{4U_o}{x_o^2} >0" .

So,it is unstable equilibrium.

For "x=+x_0,"

"-\frac{d^2U}{dx^2}=-U_o[-\frac{4}{x_o^2}+12(\frac{x}{x_o})^2(\frac{1}{x_o})^2]" "=-U_o[-\frac{4}{x_o^2}+12(\frac{x_0}{x_o})^2(\frac{1}{x_o})^2]=-\frac{8U_O}{(x_o)^2}<0"

So,it is stable equilibrium.

For "x=-x_0,"

"-\frac{d^2U}{dx^2}=-U_o[-\frac{4}{x_o^2}+12(\frac{x}{x_o})^2(\frac{1}{x_o})^2]" "=-U_o[-\frac{4}{x_o^2}+12(\frac{-x_0}{x_o})^2(\frac{1}{x_o})^2]=-\frac{8U_O}{(x_o)^2}<0"

So,it is stable equilibrium.


"(f)" "U(x) = U_o (-2 (\frac{x}{x_o})^2 + (\frac{x}{x_o})^4)"

For "x=0,"

"U(x) = U_o (-2 (\frac{x}{x_o})^2 + (\frac{x}{x_o})^4)" "=U_o (-2 (\frac{0}{x_o})^2 + (\frac{0}{x_o})^4)=0"

"\implies \frac{U}{U_o}=\frac{0}{U_o}=0"

For "x=+x_0,"

"U(x) = U_o (-2 (\frac{x}{x_o})^2 + (\frac{x}{x_o})^4)=" "U_o (-2 (\frac{+x_o}{x_o})^2 + (\frac{+x_o}{x_o})^4)=-U_o"

"\implies \frac{U}{U_o}=\frac{-U_o}{U_o}=-1"

For "x=-x_0,"

"U(x) = U_o (-2 (\frac{x}{x_o})^2 + (\frac{x}{x_o})^4)=" "U_o (-2 (\frac{-x_o}{x_o})^2 + (\frac{-x_o}{x_o})^4)=-U_o"

"\implies \frac{U}{U_o}=\frac{-U_o}{U_o}=-1"